OpenStudy (anonymous):

A shopping bag contains 6 orange juice boxes and 4 grape juice boxes. Without looking, April selects one juice box, hands it to her brother, and then takes a second juice box for herself. What is the probability that April gets an orange juice box and her brother gets a grape juice box? A. 3/5 B. 1/2 C. 4/15 D. 6/25

OpenStudy (anonymous):

I Think It's A. 3/5. ^-^

OpenStudy (anonymous):

Because 6+4=10 6/10 2x3=6 5x2=10 .... 6/10= 3/5. Right.? c:

OpenStudy (anonymous):

Start with the probability that April gets an orange juice box. Then find the probability that, after April has an orange juice box, her brother gets a grape one. Then multiply those two numbers together.